# Pressure Drop in Non-Spherical Packed Beds: Influence of Geometry and Reynolds Number

### v2

@@ Abstract
Understanding fluid flow through porous media with complex geometries is essential for improving the design and operation of packed-bed reactors. Most existing studies focus on spherical packings, having as a consequence that accurate models for irregular interstitial geometries are scarce. In this study, we numerically investigated the flow through a set of packed-bed geometries consisting of square bars stacked on top of each other and arranged in disk-shaped modules. Rotation of each module allows the generation of a variety of geometrical configurations at Reynolds numbers of up to 200 (based on the bar size). Simulations were carried out using the open-source solver OpenFOAM. Selected cases (e.g., $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\mathrm{Re}_{\mathrm{p}}=100,200$) were compared against Particle Image Velocimetry measurements. Results reveal that, based on the relative rotation angle, the realized geometries can be classified as channel-like ($\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) and lattice-like, fundamentally altering the friction factor. Furthermore, the maximum friction factor obtained in the creeping regime occurred at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, whereas in the inertial regime, this occurred at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Varying the rotation angle also affects the transition from the viscous to the inertial regime.
@@
Friction factor , Non-spherical porous medium , Packed bed , Onset of inertial regime , Non-conformal coupling
@@ 1 Introduction
However, the higher complexity of the flow field patterns induced by heterogeneous, non-spherical packings can result in large errors in velocity predictions by models describing the flows in terms of averaged velocity (Moghaddam19). Moreover, roding2017 observed that the effective diffusivity in random packings of cuboids, spheres and ellipsoids is significantly influenced by the particle shape at identical porosity. This is troubling from the perspective of unresolved flow modelling, as most models express the drag or diffusivity in terms of macroscopic parameters and would be unable to capture such differences. Standard correlations for pressure drop, such as the Ergun equation (ergun1951) or Kozeny–Carman model (carman1937), were originally developed only for spherical particles. Although these and similar drag laws have, to some extent, been successfully generalized to other geometries (e.g., liu94; duplessis1988; duplessis2008; woudberg2020), to the best of our knowledge, no studies have been conducted to quantify how the variation in packing geometry influences the induced drag at constant porosity for non-spherical packing.
Therefore, this study explores how the variation of the interstitial geometry affects the induced drag in a dense packed bed consisting of square bars. To quantify these effects, we carried out numerical simulations in a set of different geometries with the same porosity ($\phi=0.322$) at Reynolds numbers (based on bulk velocity and bar size) spanning the range between $0.1$ and $200$. Special attention was given to the influence of the geometry on the transition from viscosity- to inertia-dominated flow regimes. Moreover, to ensure the quality of the comprehensive database resulting from the 266 simulations, we carried out a detailed validation of the numerical setup against the Particle Image Velocimetry (PIV) results of christin2025. Furthermore, a thorough mesh sensitivity analysis was performed to assess the viability of the numerical approach.
@@
Figure 1: (a) The laboratory-scale packed bed of christin2025. (b) The cross-section of the packed-bed geometry formed from six modules, each rotated by $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ relative to the preceding one. (c) Schematic of the geometry of each module showing the circle enclosing the dodecagon defining the outer geometry of the slits. The streamwise direction is oriented perpendicular to the surface of the paper.
@@ 2 Geometry and Flow Conditions
The fluid moves through the column in the axial direction. The flow is characterized by a Reynolds number based on the bar width $B$ and an intrinsic average (see eq. 3) of the flow velocity through the packed bed $\langle w\rangle$:
@@
where, $\nu$ is the fluid’s kinematic viscosity.
By rotating each module by a fixed angle, different interstitial configurations and flow characteristics in the bed can be investigated. In the present study, $19$ rotation angles and, thereby, 19 geometries, are considered: $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$,$5\text{\,}\mathrm{\SIUnitSymbolDegree}$,\dotsc,$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. In order to accurately represent the macroscopic flow properties of each of packing, it is vital that a periodic geometry is considered (guibert2016; scandelli2022), otherwise a blockage effects would influence the computed values of pressure drop. For this reason, for each rotation angle, such a number of layers is chosen, which results in a periodically repeating geometry. The number of layers corresponding to each rotation angle is provided in table 1.
Figure 2 shows the simulation domains corresponding to each angle. For a small rotation angle (up to $\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), the arrangement resembles a channel-like structure, where aligned void spaces create curved flow passages. As $\alpha$ increases above $20\text{\,}\mathrm{\SIUnitSymbolDegree}$, the void spaces become interconnected and the geometry becomes more complex, with the exception of $90\text{\,}\mathrm{\SIUnitSymbolDegree}$. At $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, the geometry again exhibits a channel-like character.
@@ 3 Mathematical Model
To describe the flow at the macroscale, governing equations formulated in terms of averaged flow properties are necessary. The initial work in this field was undertaken by darcy1856, who established a linear relationship between the pressure drop and the velocity through a porous bed, under the creeping flow regime. When inertial effects become significant, the relationship between pressure gradient and velocity is no longer linear. A corrected equation accounting for fluids inertia by introducing a term proportional to a square of velocity was presented by Forchheimer1901 and is known as the Darcy–-Forchheimer equation for the pressure drop.
Although the DF model was originally postulated as empirical correlations, it has been successfully rederived by upscaling the equations valid at the pore scale (Whitaker1986; whitaker1996). Most of these derivations fundamentally rely on some form of averaging technique, such as Volume Averaging Method (VAM) (battiato2019). In the framework of VAM, an intrinsic average of a flow variable inside a porous medium is defined as
@@
where, $K_{\mathrm{eff}}=K_{\mathrm{eff}}(\mathrm{Re}_{\mathrm{p}})$ is the effective permeability of a porous material which accounts for both viscous and inertial effects, and is, therefore, a function of the Reynolds number. It can be approximated with the DF drag model formulated in terms of the isotropic permeability $K$ and the Forchheimer (or inertial) coefficient $C_{\mathrm{F}}$. The values of $K$ and $C_{\mathrm{F}}$ can be determined either experimentally, computed from simulations or estimated by using semi-empirical correlations (e.g., Lenci2022; woudberg2020; liu2024a). Alternatively, the effective permeability can be also expressed as:
@@
To compare results between different geometries, it is useful to express eq. 4 in terms on non-dimensional parameters. The pressure gradient can be normalized by the characteristic dimension of the studied geometry—taken here as BB—and the square of the superficial velocity, forming the friction factor (Dullien1975):
$f_{\mathrm{p}}=-\frac{B}{\langle w\rangle_{s}^{2}}\frac{\partial\langle p\rangle}{\partial z}.$ (6)
The permeability has the dimensions of length-squared leading to the definition of the Darcy number as $\mathrm{Da}=K/B^{2}$ and the following form of eq. 4

### v3

# Modelling of pressure drop in periodic square-bar packed beds

@@ Abstract
Understanding fluid flow through porous media with complex geometries is essential for improving the design and operation of packed-bed reactors. Most existing studies focus on spherical packings, having as a consequence that accurate models for irregular interstitial geometries are scarce. In this study, we numerically investigated the flow through a set of packed-bed geometries consisting of square bars stacked on top of each other and arranged in disk-shaped modules. Rotation of each module allows the generation of a variety of geometrical configurations at Reynolds numbers of up to 200 (based on the bar size). Simulations were carried out using the open-source solver OpenFOAM. Selected cases (e.g., $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\mathrm{Re}_{\mathrm{p}}=100,200$) were compared against Particle Image Velocimetry measurements. Results reveal that, based on the relative rotation angle, the realized geometries can be classified as channel-like ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) and lattice-like ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), fundamentally altering the friction factor. Furthermore, the maximum friction factor obtained in the creeping regime occurred at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, whereas in the inertial regime, this occurred at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. The module-equivalent diameter, based on the angle-dependent wetted surface area, collapses the friction factor onto the Ergun correlation and yields good permeability predictions for the lattice-like geometries.
@@
Pressure drop , Packed bed , Friction factor , Permeability , Tortuosity , Non-spherical porous medium
@@ 1 Introduction
However, the higher complexity of the flow field patterns induced by heterogeneous, non-spherical packings can result in large errors in velocity predictions by models describing the flows in terms of averaged velocity (Moghaddam19). Moreover, roding2017 observed that the effective diffusivity in random packings of cuboids, spheres and ellipsoids is significantly influenced by the particle shape at identical porosity. This is troubling from the perspective of unresolved flow modelling, as most models express the drag or diffusivity in terms of macroscopic parameters and would be unable to capture such differences. Standard correlations for pressure drop, such as the Ergun equation (ergun1951) or Kozeny–Carman model (carman1937), were originally developed only for spherical particles. These and similar drag laws have, to some extent, been successfully generalized to other geometries (e.g., liu94; duplessis1988; duplessis2008) and adjusted to take into account the presence of other particle shapes, either by including the sphericity of the particles as a parameter of the model (woudberg2020; Li2011EffectiveDiameter), or by constructing a multiscale model accounting for average anisotropy of the packing based on microscale simulations (tobis2008; tobis2000). That said, to the best of our knowledge, the influence on the induced drag by the variation in packing geometry at a constant macroscopic geometrical properties (e.g, porosity) remains largely unquantified.
The geometry investigated in this study is based on the modular packed-bed reactor designed by christin2025, which consists of disk-shaped modules containing square bars that can be rotated relative to one another. This allows for the study of a wide range of interstitial configurations can be studied, while keeping the porosity, bed cross-section and the shape of modules strictly constant. Although this geometry does not represent a specific industrial packing, it provides a controlled framework for studying flow through complex interstitial geometries. Crucially, the availability of detailed Particle Image Velocimetry (PIV) measurements (christin2025) enables direct validation of the numerical results, which is rarely possible for non-spherical packed beds.
Therefore, this study explores how the variation of the interstitial geometry affects the induced drag and the permeability in a dense packed bed consisting of square bars. To quantify these effects, we carried out numerical simulations in a set of different geometries with the same porosity ($\phi=0.322$) at Reynolds numbers (based on bulk velocity and bar size) spanning the range between $0.1$ and $200$. Special attention was given to the influence of the geometry on the transition from viscosity- to inertia-dominated flow regimes. Moreover, to ensure the quality of the comprehensive database resulting from the 266 simulations, we carried out a detailed validation of the numerical setup against the PIV results of christin2025. Furthermore, a thorough mesh sensitivity analysis was performed to assess the viability of the numerical approach.
@@
Figure 1: (a) The laboratory-scale packed bed of christin2025. (b) The cross-section of the packed-bed geometry (christin2025) formed from six modules, each rotated by $30\text{\,}\mathrm{\SIUnitSymbolDegree}$ relative to the preceding one. (c) Schematic of the geometry of each module showing the circle enclosing the dodecagon defining the outer geometry of the slits. The streamwise direction is oriented perpendicular to the surface of the paper.
@@ 2 Geometry and flow conditions
This geometry was designed to allow for various geometries of interstitial space: the rotation angle, height of the modules or the size of the bars can be adjusted independently. While not directly tied to a specific industrial application, the use of bars in each module enables spatio-temporally resolved optical measurements in each configuration (christin2025), unlike typically encountered, random non-spherical packing. This last property makes it valuable for studying flow characteristics via simulation approaches, as the available measurement data allows for the detailed validation at the pore scale.
By rotating each module by a fixed angle, different interstitial configurations and flow characteristics in the bed can be investigated. In the present study, $19$ rotation angles $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$,$5\text{\,}\mathrm{\SIUnitSymbolDegree}$,\dotsc,$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ and, thereby, 19 geometries, are considered. In order to accurately represent the macroscopic flow properties of each of packing, it is vital that a periodic geometry is constructed (guibert2016; scandelli2022), otherwise a blockage effects would influence the computed values of pressure drop. For this reason, for each rotation angle, such a number of layers is chosen, which results in a periodically repeating geometry. The number of layers corresponding to each rotation angle is provided in Table 1.
Figure 2 shows the simulation domains corresponding to each angle. The studied geometries can be classified by specific surface area $a_{v}$ defined as the ratio of the area of the wetted surface to the total volume of the packed bed, and the ratio of the diameter of the bed to the equivalent particle diameter $D/d_{\mathrm{eq}}$, both plotted in Fig. 3 (see Section 3 for detailed discussion of evaluating both parameters for a given geometry).
At small angles the bars in adjacent modules are nearly aligned, creating large areas of contact between the bars and yield the lowest values of $a_{v}$. As $\alpha$ increases, more surface is exposed, so that $a_{v}B$ rises to about $2.0$ near $\alpha=25^{\circ}$. For $\alpha\geq 15^{\circ}$$a_{v}B$ oscillates with $\alpha$, reflecting the periodically changing contact pattern between successive layers. Since $d_{\mathrm{eq}}\propto 1/a_{v}$ (Eqs. 17 and 8), $D/d_{\mathrm{eq}}$ follows the same trend, ranging from approximately $1.9$ at $\alpha=0^{\circ}$ to 3.13.13.33.3 for $\alpha\geq 15^{\circ}$.
For the current geometry, we picked $D/d_{\mathrm{eq}}\approx 3$ to marks the boundary between the channel-like ($D/d_{\mathrm{eq}}<3$) and lattice-like ($D/d_{\mathrm{eq}}>3$) geometries, with the transition occurring between $\alpha=10^{\circ}$ and $15^{\circ}$. Therefore, for small rotation angles ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), the arrangement resembles a channel-like structure, where aligned void spaces create curved flow passages with minimal obstruction. As $\alpha$ increases beyond $10\text{\,}\mathrm{\SIUnitSymbolDegree}$, the void spaces become increasingly interconnected and the geometry transitions to a lattice-like structure with multiply connected flow paths. The $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ configuration, with $D/d_{\mathrm{eq}}=3.23$ and $a_{v}=0.193~\mathrm{mm^{-1}}$, is classified within the lattice-like group based on these geometric parameters (see Section 5).
Figure 3: Non-dimensional specific surface area $a_{v}B$ and bed-to-particle diameter ratio $D/d_{\mathrm{eq}}$ as a function of the rotation angle $\alpha$
These values of $D/d_{\mathrm{eq}}$ place the RC2 geometry in the low bed-to-particle diameter ratio regime ($D/d<5$) (yang2016), comparable to the random packings of spheres and cylinders studied by Moghaddam19 (N=2.29N=2.296.16.1) and papkov2024 ($D/d=2.9$ and $4.8$).
The fluid moves through the column in the axial direction, with the flow conditions characterized by the pore Reynolds number The pore Reynolds number is defined as
@@
where $B$ corresponds to the bar size, $\langle w\rangle$ is the intrinsic average of the streamwise velocity, and $\nu$ is the kinematic viscosity. (In the experimental setup of (christin2025), the intrinsic average of velocity is computed directly from the volumetric flow rate forced through the packed bed.)
Depending on the application, the flows in packed beds and porous materials can be characterized by a wide range of Reynolds numbers (wood2020; Dullien1975), requiring the accurate description of both the viscous and inertial effects, to properly model the flow behaviour. Therefore, in the present paper we focus on a Range of Reynolds numbers spanning $0.1$ to $200$.
According to the flow regime classification of (Dybbs1984ANL), the transition to turbulence is typically reported at the pore-size Reynolds numbers in the range $\mathrm{Re}_{p}\approx 300$$500$ for packed beds of spheres. (wood2020), in their review of turbulent flows in porous media, confirmed that these thresholds are broadly consistent across different packing configurations, although the exact transition depends on the specific geometry, and well-defined turbulence can be established above $\mathrm{Re}_{p}\sim 750$. Since the present study is limited to $Re_{p}\leq 200$, all considered flow conditions can be expected to be laminar, with, for higher Reynolds numbers, strong inertial effects and unsteady behaviour which can exhibit intermittency.
@@ 3 Mathematical model
To describe the flow at the macroscale, governing equations formulated in terms of averaged flow properties are necessary. The initial work in this field was undertaken by darcy1856, who established a linear relationship between the pressure drop and the velocity through a porous bed, under the creeping flow regime. When inertial effects become significant, the relationship between pressure gradient and velocity is no longer linear. A corrected equation accounting for fluid inertia by introducing a term proportional to the square of the velocity was presented by Forchheimer1901 and is known as the Darcy–Forchheimer equation for the pressure drop.
Although the DF model was originally postulated as empirical correlations, it has been successfully rederived by upscaling the equations valid at the pore scale (Whitaker1986; whitaker1996). Most of these derivations fundamentally rely on some form of averaging technique, such as the Volume Averaging Method (VAM) (battiato2019). In the framework of VAM, an intrinsic average of a flow variable inside a porous medium is defined as
@@
where $K_{\mathrm{eff}}=K_{\mathrm{eff}}(\mathrm{Re}_{\mathrm{p}})$ is the effective permeability of a porous material which accounts for both viscous and inertial effects, and is, therefore, a function of the Reynolds number. It can be approximated with the DF drag model formulated in terms of the isotropic permeability $K$, which is a limit of $K_{\mathrm{eff}}$ at zero Reynolds number, and the Forchheimer (or inertial) coefficient $C_{\mathrm{F}}$. The values of $K$ and $C_{\mathrm{F}}$ can be determined experimentally, computed from simulations, or estimated using semi-empirical correlations (e.g., Lenci2022; woudberg2020; liu2024a). The effective permeability can be non-dimensionalised by selecting a characteristic length scale of the studied problem, here $B$, and forming the Darcy number $\mathrm{Da}_{\mathrm{(eff)}}=K_{\mathrm{(eff)}}/B^{2}$:
@@
It is also useful to express the pressure drop in terms of a friction factor, as commonly done in the packed-bed literature (Dullien1975). The pressure gradient can be normalized by the characteristic dimension of the studied geometry, taken here again as $B$, and the square of the superficial velocity:
$f_{\mathrm{p}}=-\frac{B}{\langle w\rangle_{s}^{2}}\frac{\partial\langle p\rangle}{\partial z},$ (6)
which combined with Eq. 5 gives:
@@
Importantly, the choice of characteristic dimension is not unique and other definitions of the friction factor have been proposed. A common approach uses the specific surface area $a_{v}=A_{p}/V_{p}$ of a particle to define the Sauter-mean diameter
$d_{\mathrm{sd}}=\frac{6}{a_{v}}=\frac{6\,V_{p}}{A_{p}}\,.$ (8)
where $V_{p}$ and $A_{p}$ are the volume and surface area of the particle, respectively. The Sauter-mean diameter does not account for particle non-sphericity. To incorporate shape effects, the sphericity $\psi$ (Wadell1935) is introduced as the ratio of the surface area of a volume-equivalent sphere to that of the actual particle:
$\psi=\frac{\pi^{1/3}\left(6\,V_{\mathrm{p}}\right)^{2/3}}{A_{\mathrm{p}}}\,.$ (9)
Li2011EffectiveDiameter proposed the equivalent particle diameter, defined as
$d_{\mathrm{eq}}=\psi\,d_{\mathrm{sd}}\,$, (10)
which reduces to $d_{\mathrm{sd}}$ for spherical particles where $\psi=1$. Using such length scale, the ergun1951 equation can be used to model the drag in packed beds of non-spherical particles (li2011), which is often expressed in the form of the following friction factor:
$f_{\mathrm{Erg}}=-\frac{\partial\langle p\rangle}{\partial z}\frac{d_{\mathrm{eq}}}{\langle w\rangle_{s}^{2}}=\frac{1-\phi}{\phi^{3}}\left(\frac{150}{\mathrm{Re}_{\mathrm{Erg}}}+1.75\right),$ (11)
with the Reynolds number defined as
$\mathrm{Re}_{\mathrm{Erg}}=\frac{\langle w\rangle d_{\mathrm{eq}}}{\nu}\frac{\phi}{1-\phi}.$ (12)
Equation 11 can be recast in the form equivalent to Eq. 4 with the Blake–Kozeny model for permeability
$K^{\mathrm{Erg}}=\frac{d_{\mathrm{eq}}^{2}}{150}\frac{\phi^{3}}{(1-\phi)^{2}}$ (13)
and the Forchheimer coefficient expressed as:
$C^{\mathrm{Erg}}_{\mathrm{F}}=\frac{1.75\,(1-\phi)\,\sqrt{K/B^{2}}}{\phi^{3}}.$ (14)
The definitions of the friction factor fundamentally depend on how the characteristic dimension is computed, as it influences both the definition of the Reynolds number and the coefficients of drag. Because the RC2 module contains interconnected bars rather than loose particles, two definitions of characteristic diameters can be considered: one based on the module geometry as a whole, and one based on a single representative bar. These are evaluated in the following section.
3.1 Geometric parameters for the RC2 configuration
In the present work, the specific surface area is primarily defined as
$a_{v}=\frac{A_{\mathrm{wetted}}}{V_{\mathrm{total}}}\,$, (15)
where $A_{\mathrm{wetted}}$ is the total wetted surface area of the module and $V_{\mathrm{total}}$ is the total module volume (comprising both fluid and solid regions). Because the bar-to-bar contact regions change with the rotation angle $\alpha$, so does the wetted surface exposed to the fluid, and consequently both $a_{v}$ and the bed-to-particle diameter ratio $D/d_{\mathrm{eq}}$ vary with $\alpha$ (Fig. 3). The equivalent diameter $d_{\mathrm{eq}}$ is derived from the module as a whole. The module-based sphericity is
$\psi_{\mathrm{eq}}=\frac{\pi^{1/3}\,\bigl(6\,V_{\mathrm{p}}\bigr)^{2/3}}{A_{\mathrm{wetted}}}\,$, (16)
The Sauter-mean diameter follows directly from Eq. 8 and likewise depends on $\alpha$. The module-based equivalent diameter then follows from Eqs. 10 and 8 as
$d_{\mathrm{eq}}=\psi_{\mathrm{eq}}\,d_{\mathrm{sd}}^{\mathrm{eq}}\,.$ (17)
One can also propose a characteristic length scale by establishing a representative particle and use it to compute geometric properties: based on a single representative square bar, we define the single-bar diameter $d_{\mathrm{sb}}$. The simulation domain per module contains four square bars arranged in two symmetrically positioned pairs that share the same cross-section ($10\times 10\;\mathrm{mm^{2}}$) but differ in length ($L_{1}=62\;\mathrm{mm}$ and $L_{2}=46\;\mathrm{mm}$). A representative bar is therefore defined with the averaged length $L_{\mathrm{avg}}=(L_{1}+L_{2})/2=54\;\mathrm{mm}$, giving $V_{\mathrm{sb}}=10\times 10\times 54=5400\;\mathrm{mm^{3}}$ and $A_{\mathrm{sb}}=2\,(10\times 10)+4\,(10\times 54)=2360\;\mathrm{mm^{2}}$. The single-bar sphericity is
$\psi_{\mathrm{sb}}=\frac{\pi^{1/3}\,\bigl(6\,V_{\mathrm{sb}}\bigr)^{2/3}}{A_{\mathrm{sb}}}\,$, (18)
where $V_{\mathrm{sb}}$ and $A_{\mathrm{sb}}$ are the volume and surface area of the representative square bar, respectively. For the representative bar described above, $\psi_{\mathrm{sb}}=0.63$. The Sauter-mean diameter for a single bar is
$d_{\mathrm{sd}}^{\mathrm{sb}}=\frac{6\,V_{\mathrm{sb}}}{A_{\mathrm{sb}}}\,.$ (19)
The corresponding single-bar equivalent diameter reads
$d_{\mathrm{sb}}=\psi_{\mathrm{sb}}\,d_{\mathrm{sd}}^{\mathrm{sb}}\,.$ (20)
Note that $d_{\mathrm{sb}}$ is constant for all $\alpha$, since both $\psi_{\mathrm{sb}}$ and $d_{\mathrm{sd}}^{\mathrm{sb}}$ are determined by the fixed bar geometry, yielding $d_{\mathrm{sb}}=8.65\;\mathrm{mm}$. The Blake–Kozeny permeability estimate (Eq. 13) can be evaluated with each characteristic diameter:
$K_{\mathrm{eq}}=\frac{d_{\mathrm{eq}}^{2}}{150}\,\frac{\phi^{3}}{(1-\phi)^{2}}\,,\qquad K_{\mathrm{sb}}=\frac{d_{\mathrm{sb}}^{2}}{150}\,\frac{\phi^{3}}{(1-\phi)^{2}}\,.$ (21)
For both diameter definitions, the Ergun-type Reynolds number (Eq. 12) is formed with the respective equivalent diameter:
$\mathrm{Re}_{\mathrm{Erg}}^{d_{\mathrm{eq}}}=\frac{\langle w\rangle\,d_{\mathrm{eq}}}{\nu}\,\frac{\phi}{1-\phi}\,,\qquad\mathrm{Re}_{\mathrm{Erg}}^{d_{\mathrm{sb}}}=\frac{\langle w\rangle\,d_{\mathrm{sb}}}{\nu}\,\frac{\phi}{1-\phi}\,.$ (22)
3.2 Tortuosity-based permeability model
The Blake–Kozeny and Kozeny–Carman models can be both expressed in a general sense as
$K=\frac{\phi^{3}}{c_{\mathrm{KC}}\,\tau^{2}\,a_{v}^{2}}\,$, (23)
where $c_{\mathrm{KC}}$ is the Kozeny constant (for a cylindrical capillary $c_{\mathrm{KC}}=2$) and $\tau$ is the Hydraulic tortuosity which quantifies the elongation of fluid paths relative to the straight-line distance through the medium. For spheres, $\tau$ is typically fixed at $\tau=\sqrt{2}$ (clennell1997; carman1937) leading to the Blake–Kozeny model.
Naturally, there is no certainty that a model formulated with a constant value of $\tau$ will be valid across many geometries and estimating permeability by the use of Eq. 23 may yield improved results. Different definitions of tortuosity have been proposed in the literature (bear1972; clennell1997; matyka2008; duda2011). Following duda2011, the upper bound for the hydraulic tortuosity can be computed directly from the velocity field as the ratio of the volume-averaged velocity magnitude to the volume-averaged axial velocity component in the void space:
$\tau_{\mathrm{eff}}=\frac{\langle|\bm{u}|\rangle}{\langle w\rangle}.$ (24)
At the larger Reynolds numbers, the upper bound for the hydraulic tortuosity depends on $\mathrm{Re}_{\mathrm{p}}$ through the rearrangement of the flow field by inertial effects and is denoted $\tau_{\mathrm{eff}}$. In the Darcy limit, it reduces to
$\tau=\lim_{\mathrm{Re}_{\mathrm{p}}\to 0}\tau_{\mathrm{eff}}\,$, (25)
which is a purely geometric property of the void space.